In the year 2009,the average annual cost to employees for employer family health coverage was approximately ​$3500.In 2014,that amount was approximately ​$4500.
​(a) Write a linear equation expressing the average annual cost to employees y in terms of the number of years x after 2000.
​(b) What was the average annual cost in​ 2014?
​(c) If the trend​ continues, what will be the first full year in which costs exceed ​$7000?

Answer by Boreal(15235)

2009=9 and 2014=14
we have ordered pairs (9, 3500) and (14, 4500)
that slope is 1000/5 or 200
point slope formula is y-y1=m(x-x1), m slope (x1, y1) point
y-3500=200(x-9)
y=200x-1800
y=200x+1700
in 2014, the cost was 200(14)+1700 or $4500, which we knew
7000=200x+1700
200x=5300
x=26.5
That means in 2026, the cost is $6900 and in 2027 it is $7100, so 2027.